The velocity of body at time t=0 is 10√2 m/s  in North East direction as it is moving with an acceleration of 2 m/s2 direcr towards south. The magnitude and direction of body after 5 sec will be..

Grade:12th pass

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

Velocity = 10i + 10j

acceleration = -2j
time 5 sec

Please specify magnitude and direction of WHAT (SPEED / DISTANCE / ACCELERATION)

use v=u+at; s=ut+(0.5)at2; v2=u2 +2as;

Regards